Can you help me find the answer for this question I have tried using the probability and get confused when there are two subjects the question is asking. What is the probability of selecting a student who liked Chemistry or Biology?
Major
Gender Biology Anatomy Chemistry Total
Male 105 140 47 292
Female 103 25 52 180
Total 208 165 99 472

Found 2 solutions by sudhanshu_kmr, stanbon:
Answer by sudhanshu_kmr(1152)


probability = (208 +99)/ 472

= 307/472

Answer by stanbon(75887)

What is the probability of selecting a student who liked Chemistry or Biology?
Major
Gender Biology Anatomy Chemistry Total
Male 105 140 47 292
Female 103 25 52 180
Total 208 165 99 472
-----------
Ans: P(chem or biol) = P(chem) + P(biol) = 47/472 + 105/472 = 152/472
---------------
Cheers,
Stan H.